# Exploring specialization and sensitivity of convolutional neural networks in the context of simultaneous image augmentations

### v1

@@ 1 Introduction
In this work, we have employed the sensitivity analysis based on Sobol indices and Shapley values for highlighting the network’s components that give the most prominent feedback in response to the input augmentations. As a result, we introduce the framework of selecting and studying the units sensible to the input data augmentations (Figure 1). The variables incorporated in the analysis are augmentation-related, augmentation-permuting, partition- and class-selecting ones. Computed values are plotted as sensitivity maps, which were investigated manually and used as features for correlational and linear discriminant analysis. This analysis enabled us to identify internal neural processing tendencies layerwise and on a single unit scale, validated by damaging the convolutional checkpoints with a guided masking prediction as an independent procedure that matches sensitivity values with corresponding augmentations. In addition, we performed the single-class sensitivity analysis as an effort to understand the sensitivity of the whole network on the basis of the classifying layer sensitivity. As evidence that these findings are generalizable, we seek for prediction bias - more frequent prediction of the classes identified by the classifying layer as sensible to an augmentation - generated by the guided masking of convolutional checkpoints. Finally, the framework was adapted to analyze the network segments. In the example of AlexNet [29], VGG11 [30], and ResNet18 [31] models, we demonstrate the alignment among different analysis blocks and consistency with the external knowledge, including patterns observed in biological circuits. We also recognize our findings as potentially transferable for studying biological neural networks in complex environments.
@@ 2.5 Experimental setup
Computed sensitivity values (Figures S1–S6) were visually inspected and analyzed via extracting correlation and prediction patterns. The correlation patterns were evaluated by calculating unit-wise Spearman correlations between SA variables (Figures S7.2(a)–(e)). Both augmentation sets were simultaneously considered in order to relate them. Units (output channels) of the fixed checkpoint were treated as observations. In the case of convolutional checkpoints, the results were spatially averaged. In addition to resulting correlation matrices, spatial correlation maps were generated to reflect the possible relationships between sensitivity to augmentations and coefficients of variations as a measure of unit activity (Figures S7.3(a)–(c)).
Additional exploratory analysis was performed by extracting prediction patterns from confusion matrices (Figures S7.4(a)–(c)). A convolutional neuron maps a multi-channelled image to a single-channelled one with sensitivity values to different SA variables assigned to each point. They were employed as features for predicting the SA variables. Resulting confusion matrices highlight high-level relationships between SA variables regarding linear separability patterns. As a separation method, we used cross-validated linear discriminant analysis (LDA): all units from a fixed convolutional checkpoint were split into training and validation sets by applying the K-fold scheme () 100 times. Predictions were measured on the validation parts, and corresponding confusion matrices were computed as averages.
@@
Sensitivity analysis marks feature maps with a variability level of activations in response to an applied augmentation. The resulting ‘‘markup’’ may be employed as a mask outlining more and less sensitive areas, which makes it possible to perform guided editing of the activations. Such perturbations break the communication system established among neurons from neighbouring layers. It can be compared to manual interference with synaptic chemistry by blocking (or activating) sensible receptors located in the synaptic terminal, changing the chemical context and leaving the location of synapses and neurons unchanged. It is a variant of reversible brain damage, similar to the temporary effect caused by injecting a specific drug. The same idea stands behind the Dropout regularizing method, with the distinction that the latter completely zeroes out a subset of the output during the learning procedure.
@@
While investigating the sensitivity values, we observed them as skewed and non-uniform. Thus, we considered the problem of detecting the single-class sensitivity. The first-order Sobol indices and Shapley values were engaged to address this problem. Total Sobol indices were excluded from the consideration due to their weaker specificity. The candidate list may be ranked by a simple sorting in descending order of the sensitivity values. However, where to put a threshold separating the non-sensible classes is not apparent. Our work employed a top-5 approach as a rule of thumb (Tables 2, 3). Using the same masking approach, we created a new pattern based on top-5 predictions to filter the results. Predictions and sensible classes were matched between each other by the Jaccard index, which increases as predictions become more biased towards the sensible classes (Tables 4, S7.6(a)–(f)).
@@
In the last experimental series, we feed the data directly to the standalone segments of neighbouring checkpoints. As the number of input channels does not match the input data, we represented original images in HSV format (see Figure 2) and separately repeated each of the resulting channels to fit the corresponding segment (single-channel checkpoint). Then, we extracted the pixel-wise correlation patterns for different components of HSV decomposition with the original values from the first experimental series on the same manually generated subset of units recognized as the most sensible to at least one of the augmentations (Figures S7.7(a)–(d)). These experiments were conducted only on AlexNet and ResNet18.
@@ 3.1.1 SA: total variances
Before studying the sensitivity values, indicating proportional variable contribution, we explore variances and coefficients of variation. As shown in Figures S7.1.1–S7.1.3, (a)–(b), distribution patterns for variances are similar for both sampling schemes and vary depending on the augmentation set. Since zero value of activations means zero contribution during inference, we considered the coefficient of variation (CoV) as a ratio of standard deviation to the absolute value of means to compare units directly (Figures S7.1.1–S7.1.3, (c)–(d)). They tend to be more uniform and more deviated from zero on the whole, suggesting that most units are involved in inference. No connection can be drawn between the units most sensitive to a single augmentation and those with high variances or CoVs (Figures S7.1.1–S7.1.3, Figures S1–S6, Figure 3).
@@ 3.1.2 SA: neuronal activation patterns
Standalone constant sharpness and horizontal flip augmentations weakly influence the prediction (Table 1) but demonstrate different sensitivity patterns: there is a low sensitivity for sharpness adjustment (Shapley values, first-order Sobol indices) at convolutional checkpoints. In contrast, the horizontal flip has structured sensitivity patterns (Shapley values).
Rolling generates either border bar patterns (first-order Sobol) or more uniform global sensitivity involving interactions (Shapley values). Local patterns appear for erasing and elliptic local blur augmentations (first-order Sobol, Shapley). Erasing affects both the predictions and activations severely, i.e., networks more easily compensate for the elliptic blurring as more local transform which saves more information.
@@ 3.1.3 Correlation patterns
Correlation matrices between SA variables are plotted in Figures S7.2(a)–(e). Correlations within the augmentation set tend towards a common level as checkpoints progress deeper, clustering variables together and smoothing their impact. Grayscale and saturation/hue have a strong correlation by means of the first-order Sobol indices and Shapley values. For the beginning checkpoints, it correlates more with saturation, giving way to a correlation with hue in the deeper layers. In both cases, these correlations decay to zero as checkpoints move closer to the avgpool (however, Alexnet breaks the pattern exactly at this checkpoint). Hue and saturation variables are also correlated, but their correlation decays rapidly to zero or their cluster’s common level. Other notable positive correlations are observed for pairs from and brightness–grayscale pair, in terms of Shapley values. Total Sobol indices distinguish the erasing transform as significantly spaced away from other augmentations, forming the distinct cluster for several checkpoints closer to the beginning of a network. The trend is slightly less expressed for a hue variable and even less for grayscale and saturation.
@@ 3.1.4 Spatial correlation maps
Spatial correlation maps relating the SA variables and CoVs can be found in Supplementary materials as Figures S7.3(a)–(c). The class variable has near-zero correlations for total Sobol indices and positive or near-zero (AlexNet/VGG11, augmentation set 2) for the first-order Sobol indices. The final-layer sensitivity by means of Sobol indices shares the inhibition pattern for single augmentations, i.e., the activations loose augmentation-explained variability. Correlation patterns based on Shapley values for the last layer are more diverse, and the common trend for all considered networks is the same near-zero or weak positive correlation for the class variable and weak negative correlation for the partition variable. For convolutional checkpoints, there are complete correlation maps, which may be viewed as spatial on-off patterns similar to ones in biological networks: positive correlation areas reflect the level of aligned matching between sensitivity and activation corresponding to excitation patterns, and negative correlations can be related to inhibition patterns. And similarly to biological networks [47, 48], the corresponding receptive fields demonstrate center-surrounding and distributed structure. These spatial patterns depend on network, checkpoint, and SA variable, as well as on sensitivity values. Distributed patterns are minor for total Sobol indices, while the first-order Sobol indices and Shapley values exhibit both. As in biological neurons, the on-off fluctuations exist [49]: for a fixed network, the class or partition spatial correlation maps are similar in general but locally different depending on the augmentation set.
@@ 3.1.5 Linear discriminant analysis
Confusion matrices derived from cross-validated LDA are presented in Figures S7.4(a)–(c). These results are detailed in B; here, we summarize the common tendencies. SA variables are well separated at the first checkpoint for all values and networks. This separability decays while moving deeper, which is agreed with previously identified correlational patterns. Strong separation of erasing transform lasts longer again, even in the deepest checkpoints. The possible explanation for this pattern is that it causes severe distortions which are difficult to compensate for and easy to detect even with a linear classifier. Also, the closer to the classifying layer a checkpoint is located, the more separable the class variable becomes - there are either U-shaped or nearly parallel trends, along with poorer separability of certain augmentation variables. That may stem from the cancellation of augmentation impact and retargeting of neurons to the class variable as a network complying with the inference’s goal and learning higher-level abstractions.
@@ 3.2.1 Correlation patterns
Figure 4 presents the matching results for input augmentations and guided activation masking. Certain pairs for Shapley values and first-order Sobol indices are the most selective: Gaussian blur, grayscale, erasing, and hue transforms are the most diverging when masking and augmentation variables coincide, resulting in lower diagonal values. It supports the matching between computed sensitivity values and real network sensitivity. Total Sobol indices break this pattern for most cases (except grayscale and a few other partial entries), presumably due to their structure incorporating higher-level interactions among SA variables.
@@
On the first augmentation set, the erasing transform separates well for Shapley values and worse for Sobol indices (AlexNet, erasing-permutation pair dominates erasing-erasing; ResNet18 - all values are high); for total Sobol indices, entries deviate significantly for all pairs. The grayscaling is separable in the case of AlexNet and VGG11 for all sensitivity values, on ResNet18 all its entries are close to one for Shapley values, and for first-order and total Sobol indices all these pairs are diverging. The Gaussian blur’s matching is impaired on VGG11 for Shapley values and all networks for total Sobol indices. The rolling is discriminated for Sobol indices, on AlexNet - also for Shapley values.
@@
On the second augmentation set, AlexNet’s and VGG11’s results are comparable again: both networks signalled the selective sensitivity to the elliptic local blurring and rotation for Sobol indices. The horizontal flipping resulted in close to one correlation for all sensitivity values and networks. The rest colour adjustments are highly cohesive. The brightness is separable well in many cases (even for total Sobol indices for VGG11 and ResNet18), except for a saturation pairing. Another well-discriminative transform is the hue adjustment (also separable for ResNet18’s total Sobol indices), which is also confused with saturation for Sobol and Shapley values. The saturation itself is mostly entangled with hue, and also with contrast as for ResNet18. However, the contrast is well-matching its masking for all sensitivity values on ResNet18, and tangled with saturation on AlexNet and VGG11 for Shapley values, with all colour pairings for first-order Sobol indices.
@@
In addition, for erasing, hue, contrast, Gaussian blur and grayscale augmentations masking with the other variables affects predictions significantly, indicating the possible instabilities related to these augmentations. This view is supported by the baseline measurements given in Table 1, as well as the patterns described in the previous subsection. It may also indicate the intersections among sensitivity maps, and the more local it is, the weaker the prediction pattern affected (for instance, horizontal flipping, local elliptic blur; erasing is not so localized, especially in deeper checkpoints). Although sharpness adjustment by a constant factor has little effect on regular predictions, it is a global augmentation; hence, other variables influence its prediction pattern. Interestingly, it is affected mainly by a Gaussian blur variable, which operates with the same abstract property (sharpness).

### v2

@@ 1 Introduction
Compared to certain existing explainability methods, we propose to focus on how the network propagates input distortions themselves instead of using them to build attribution maps for specific image samples (as, for instance, in [16, 28, 29, 30]). Although there are known studies that also explore the influence of various types of distortion [31, 32, 33, 34, 35], to the best of our knowledge, simultaneous input augmentations have not been presented yet. We see our main contribution as addressing this issue and proposing a framework that is able to account for potential dependencies among injected distortions, as well as for other variables affecting input data (such as class or data partition). As dependencies rise even from the non-permutability of the sequential augmentations, their order is another crucial factor incorporated into our analysis.
In this work, we have employed the sensitivity analysis based on Sobol indices and Shapley values for highlighting the network’s components that give the most prominent feedback in response to the input augmentations. As a result, we introduce the framework of selecting and studying the units sensitive to the input data augmentations (Figure 1). The variables incorporated in the analysis are augmentation-related, augmentation-permuting, partition- and class-selecting ones. Computed values are plotted as sensitivity maps, which were investigated manually and used as features for correlational and linear discriminant analysis. This analysis enabled us to identify internal neural processing tendencies layerwise and on a single unit scale, validated by damaging the convolutional checkpoints with a guided masking prediction as an independent procedure that matches sensitivity values with corresponding augmentations. In addition, we performed the single-class sensitivity analysis as an effort to understand the sensitivity of the whole network on the basis of the classifying layer sensitivity. As evidence that these findings are generalizable, we seek for prediction bias - more frequent prediction of the classes identified by the classifying layer as sensitive to an augmentation - generated by the guided masking of convolutional checkpoints. Finally, the framework was adapted to analyze the network segments. In the example of AlexNet [36], VGG11 [37], and ResNet18 [38] models, we demonstrate the alignment among different analysis blocks and consistency with the external knowledge, including patterns observed in biological circuits. We also recognize our findings as potentially transferable for studying biological neural networks in complex environments.
@@ 2.5 Experimental setup
Computed sensitivity values (Figures S1–S6) were visually inspected and analyzed via extracting correlation and prediction patterns. The correlation patterns were evaluated by calculating unit-wise Spearman correlations between SA variables (Figures S7.2). Both augmentation sets were simultaneously considered in order to relate them. Units (output channels) of the fixed checkpoint were treated as observations. In the case of convolutional checkpoints, the results were spatially averaged. In addition to resulting correlation matrices, spatial correlation maps were generated to reflect the possible relationships between sensitivity to augmentations and coefficients of variations as a measure of unit activity (Figures S7.3).
Additional exploratory analysis was performed by extracting prediction patterns from confusion matrices (Figures S7.4). A convolutional neuron maps a multi-channelled image to a single-channelled one with sensitivity values to different SA variables assigned to each point. They were employed as features for predicting the SA variables. Resulting confusion matrices highlight high-level relationships between SA variables regarding linear separability patterns. As a separation method, we used cross-validated linear discriminant analysis (LDA): all units from a fixed convolutional checkpoint were split into training and validation sets by applying the K-fold scheme () 100 times. Predictions were measured on the validation parts, and corresponding confusion matrices were computed as averages.
@@
Sensitivity analysis assigns a variability level of activations in response to an applied augmentation to feature maps. The resulting ‘‘markup’’ may be employed as a mask outlining more and less sensitive areas, enabling guided editing of the activations. Such perturbations break the communication pathways established among neurons from neighbouring layers. It resembles a manual blocking (or activating) sensitive receptors located in the synaptic terminal, changing the chemical context while leaving the location of synapses and neurons unchanged, i.e., a variant of reversible brain damage, similar to the temporary effect caused by injecting a specific drug. The same idea stands behind the Dropout regularizing method, with the distinction that the latter completely zeroes out a subset of the output during the learning procedure.
Besides, there is a masking-based technique to estimate the sensitivity of occlusion attribution (e.g., [29]), where masking is applied as a single transformation to identify specific areas of an input exerting the strongest influence on the prediction. Recall that in our work we directly use the intermediate changes of activations derived from different combinations of various input augmentations. Sensitivity values allows one to assess the contibution of every SA variable to the activations’ variance. But how specific are the resulting sensitivity maps? Figure 3 demonstrates that sensitivity to several SA variables yields a compensatory effect for others. Moreover, there are intersections of augmentation actions, e.g., modifying the high-level feature like luminosity or sharpness. Therefore, the objective of the masked activation analysis comprises an integral specificity assessment of the sensitivity maps produced during simultaneous augmenting setup through matching them with single augmentations. Activation editing should not be mistaken with input augmentations: the latters are performed independently on any sensitivity maps, and activations within selected convolutional checkpoints are to be masked according to estimated sensitivity maps.
@@
While investigating the sensitivity values, we observed them as skewed and non-uniform. Thus, we considered a problem of detecting the single-class sensitivity. The first-order Sobol indices and Shapley values were engaged to address this problem. Total Sobol indices were excluded from the consideration due to their weaker specificity. The candidate list may be ranked by a simple sorting in descending order of the sensitivity values. However, where to put a threshold separating the non-sensitive classes is not apparent. Our work employed a top-5 approach as a rule of thumb (Tables 2, 3). Using the same masking approach, we created a new pattern based on top-5 predictions to filter the results. Predictions and sensitive classes were matched between each other by the Jaccard index, which increases as predictions become more biased towards the sensitive classes (Tables 5, S7.6).
@@
In the last experimental series, we feed the data directly to the standalone segments of neighbouring checkpoints. As the number of input channels does not match the input data, we represented original images in HSV format (see Figure 2) and separately repeated each of the resulting channels to fit the corresponding segment (single-channel checkpoint). Then, we extracted the pixel-wise correlation patterns for different components of HSV decomposition with the original values from the first experimental series on the same manually generated subset of units recognized as the most sensitive to at least one of the augmentations (Figures S7.7). These experiments were conducted only on AlexNet and ResNet18.
@@ 3.1.1 SA: total variances
Before studying the sensitivity values, indicating proportional variable contribution, we explore variances and coefficients of variation. As shown in Figures S7.1.1–S7.1.5, (a)–(b), distribution patterns for variances are similar for both sampling schemes and vary depending on the augmentation set. Since zero value of activations means zero contribution during inference, we considered the coefficient of variation (CoV) as a ratio of standard deviation to the absolute value of means to compare units directly (Figures S7.1.1–S7.1.5, (c)–(d)). They tend to be more uniform and more deviated from zero on the whole, suggesting that most units are involved in inference. No connection can be drawn between the units most sensitive to a single augmentation and those with high variances or CoVs (Figures S7.1.1–S7.1.5, Figures S1–S6, Figure 3).
The exception is AlexNet trained on the Places 365 dataset, where many neurons of the second, third and partially first checkpoints are inactive (Figures S7.1.4). As the network was trained from a scratch [54] (see also a link from Data availablility), an inproper weights initialization could cause the outcome. This issue impacts all further steps of our analysis (hereafter, we omit the remark), providing additional data on the framework’s robustness. ResNet18 did not surface the problem, presumably due to its skip-connection feature. Apart from that, we observed the similar tendencies as for ILSVRC.
@@ 3.1.2 SA: neuronal activation patterns
Standalone constant sharpness and horizontal flip augmentations weakly influence the prediction (Table 1) but demonstrate different sensitivity patterns during simultaneous augmenting setup: there is a low sensitivity for sharpness adjustment (Shapley values, first-order Sobol indices) at convolutional checkpoints. In contrast, the horizontal flip has structured sensitivity patterns (Shapley values).
Rolling generates either border bar patterns (first-order Sobol) or more uniform global sensitivity involving interactions (Shapley values). Local patterns appear for erasing and elliptic local blur augmentations (first-order Sobol, Shapley). Erasing affects both the predictions and activations severely, i.e., networks more easily compensate for the elliptic blurring as more local transform which preserves more information.
@@ 3.1.3 Correlation patterns
Correlation matrices between SA variables are plotted in Figures S7.2.1 (a)–(e), S7.2.2 (a)–(c). Correlations within the augmentation set tend towards a common level as checkpoints progress deeper, clustering variables together and smoothing their impact. Grayscale and saturation/hue have a strong correlation by means of the first-order Sobol indices and Shapley values. For the beginning checkpoints, it tends to correlate more with saturation, giving way to a correlation with hue in the deeper layers. In both cases, these correlations decay to zero as checkpoints move closer to the avgpool (however, Alexnet breaks the pattern exactly at this checkpoint). Hue and saturation variables are also correlated, and their correlation decays to zero or their cluster’s common level. Other notable positive correlations are observed for pairs from , in terms of Shapley values. Total Sobol indices distinguish the erasing transform as significantly spaced away from other augmentations, forming the distinct cluster for several checkpoints closer to the beginning of a network. A similar trend to constitute a separate cluster is slightly less pronounced for a hue variable and even less prominent for grayscale and saturation.
@@ 3.1.4 Spatial correlation maps
Spatial correlation maps relating the SA variables and CoVs can be found in Supplementary materials as Figures S7.3.1 (a)–(c), S7.3.2 (a)–(b). The class variable has near-zero correlations for total Sobol indices measured at the final (classifying) checkpoint. In addition, for single augmentations the final-layer sensitivity shares either the inhibition pattern or near-zero correlations, suggesting that the activations loose augmentation-explained variability there. Correlation patterns based on Shapley values for the last layer are more diverse than ones for Sobol indices, and the common trend for all considered networks is the same near-zero or weak positive correlation for the class variable and weak negative correlation for the partition variable. For convolutional checkpoints, there are complete correlation maps, which may be viewed as spatial on-off patterns similar to ones in biological networks: positive correlation areas reflect the level of aligned matching between sensitivity and activation corresponding to excitation patterns, and negative correlations can be related to inhibition patterns. And similarly to biological networks [55, 56], the corresponding receptive fields demonstrate both center-surrounding and distributed structures. These spatial patterns depend on network, dataset, checkpoint, and SA variable, as well as on sensitivity values. Distributed patterns are minor for total Sobol indices, while the first-order Sobol indices and Shapley values exhibit both. As in biological neurons, the on-off fluctuations exist [57]: for a fixed network, the class or partition spatial correlation maps are similar in general but locally different depending on the augmentation set.
@@ 3.1.5 Linear discriminant analysis
Confusion matrices derived from cross-validated LDA are presented in Figures S7.4.1 (a)–(c), S7.4.2 (a)–(b). These results are detailed in B; here, we summarize the common tendencies. SA variables are well separated at the first checkpoint for all values and networks except AlexNet trained on the Places 365 dataset (see remark in 3.1.1). This separability decays while moving deeper, which is agreed with previously identified correlational patterns. Strong separation of erasing transform lasts longer again, even in the deepest checkpoints. The possible explanation for this pattern is that it causes severe distortions which are difficult to compensate for and easy to detect even with a linear classifier. Also, the closer to the classifying layer a checkpoint is located, the more separable the class variable becomes - there are either U-shaped or nearly parallel trends, along with poorer separability of certain augmentation variables. That may stem from the cancellation of augmentation impact and retargeting of neurons to the class variable as a network complying with the inference’s goal and learning higher-level abstractions.
@@ 3.2.1 Correlation patterns
Figure 4 presents the matching results for input augmentations and guided activation masking. Certain pairs for Shapley values and first-order Sobol indices are the most selective: Gaussian blur, grayscale, erasing, and hue transforms are the most diverging when masking and augmentation variables coincide, resulting in lower diagonal values. It supports the matching between computed sensitivity values and real network sensitivity. Total Sobol indices break this pattern in most cases (except grayscale and a few other partial entries), presumably due to their structure incorporating higher-level interactions among SA variables.
@@
On the first augmentation set, the erasing transform separates well for Shapley values and for Sobol indices (exceptions: AlexNet, erasing-permutation pair dominates erasing-erasing on the ImageNet dataset; ResNet18 - all values are high, but still separable on the Places 365); for total Sobol indices, entries deviate significantly on all pairs. The grayscaling is separable in the case of AlexNet and VGG11 for all sensitivity values; ResNet18’s entries are either close to one (ImageNet, Shapley) or diverging. The Gaussian blur’s matching is rarely impaired (VGG11, Shapley values; all networks, total Sobol indices). The rolling is well-discriminated for Sobol indices (except for ResNet18, Places 365) and for AlexNet’s Shapley values (ImageNet).
@@
On the second augmentation set, AlexNet’s and VGG11’s results are aligned on the ImageNet dataset: both networks showed selective sensitivity to elliptical local blurring and rotation for Sobol indices. The horizontal flipping resulted in a correlation coefficient of nearly one for all sensitivity values, networks, and datasets. The rest, colour adjustments, are highly cohesive. The brightness is separable in many cases (even for total Sobol indices, VGG11 and ResNet18, ImageNet), but susceptible to confounding with other SA variables in the case of ResNet18 (Shapley values, the first-order Sobol indices). The hue adjustment is well-paired with the same SA variable masking, although partially overlapped with the saturation variable. The saturation itself is mostly entangled with hue and slightly with contrast (ResNet18, ImageNet; AlexNet, Places 365). The contrast transformation matches well with its masking for all ResNet18’s sensitivity values (with the only exception, Sobol indices on the Places 365), though it is highly affected by other SA variables (especially other colour adjustments) and tangled with saturation for AlexNet’s and VGG11’s Shapley values.
@@
In addition, for erasing, hue, contrast, Gaussian blur, and grayscale augmentations, masking with the other variables significantly affects predictions, indicating the possible instabilities related to these augmentations. This view is supported by the baseline measurements given in Table 1, as well as the patterns described in the previous subsection. It may also indicate the intersections among sensitivity maps, and the more local it is, the weaker the prediction pattern affected (for instance, horizontal flipping, local elliptic blur; erasing is not so localized, especially in deeper checkpoints). Although sharpness adjustment by a constant factor has little effect on regular predictions, it is a global augmentation; hence, other variables influence its prediction pattern. Interestingly, on ImageNet, it is affected mainly by a Gaussian blur variable, which operates with the same abstract property (sharpness). As for Places 365, this pattern is reversed (ResNet18), and the blurred input tends to be affected by the sharpness variable.